# LLM Agents for Education: Advances and Applications

### v1

@@ 5. Challenges & Future Directions
In this section, we discuss key challenges that must be addressed to ensure the effectiveness, reliability, and ethical deployment of LLM agents for education. We focus on three critical areas: privacy, bias, and fairness concerns; hallucination and overreliance; and integration with existing educational ecosystems. For each challenge, we outline potential research directions to enhance the robustness and applicability of LLM agents in learning environments.
5.1. Privacy, Bias and Fairness Considerations
The integration of LLM agents into educational applications offers transformative potential but also raises significant ethical concerns, particularly regarding privacy and bias.
@@ 5.1. Privacy, Bias and Fairness Considerations
LLM agents process vast datasets, often containing sensitive personal information, leading to potential privacy risks. Studies highlight low technological readiness and insufficient privacy measures in educational contexts (Yan et al., 2024a). Emerging research (He et al., 2024b; Gan et al., 2024; Zhang et al., 2024d; Hua et al., 2024; Huo et al., 2025; Chen et al., 2025a) underscores new privacy and security concerns, emphasizing the need for stronger data protection mechanisms (Huang, 2023; Ismail, 2025; Khan, 2024). Additionally, bias in LLMs remains a pressing concern, as models trained on large datasets can inadvertently reinforce stereotypes and disparities, affecting educational fairness. Recent work calls for bias mitigation strategies to promote equitable learning experiences (Adewumi et al., 2024; Aird et al., 2024; Mehrotra et al., 2024). Addressing these biases is essential to ensuring inclusive, unbiased educational outcomes.
@@
To overcome the above issues, a number of future directions can be explored: (i) Privacy-preserving memory management: develop context-aware memory mechanisms that allow agents to retain useful learning progress while forgetting sensitive user data when necessary. (ii) Bias detection and mitigation: develop automated fairness-checking models that evaluate real-time content generated by LLM agents to detect biased explanations, language, or examples. (iii) Culturally adaptive LLM Agents for global education: train multilingual, culturally aware educational agents that dynamically adjust explanations based on regional learning norms, historical perspectives, and diverse curricula.
5.2. Hallucination and Overreliance
The phenomenon of ”hallucinations” (Zhang et al., 2023), where LLMs generate plausible-sounding but incorrect or nonsensical information, poses challenges to their reliability in educational contexts.
In educational settings, hallucinations can mislead learners by presenting plausible but inaccurate information as factual, leading to misconceptions (da Silva et al., 2024; Jho, 2024; Ho et al., 2024). For example, AI-generated content may fabricate historical events or scientific data, which students may unknowingly accept as true. This issue is particularly concerning given the authoritative tone of AI-generated responses, making errors harder to detect and correct. Also, recent researches (Milano et al., 2023; Krupp et al., 2024; Adewumi et al., 2023) show that overreliance on AI-generated content may hinder genuine skill acquisition and impede in-depth learning.
@@ 5.2. Hallucination and Overreliance
Some directions can be explored to mitigate hallucinations in LLM agents for education: (i) Self-correcting AI tutors: develop LLM agents with self-reflection capabilities, where models review, verify, and refine their own generated content before presenting it to students. (ii) Hybrid Human-AI feedback loops for educational content verification: develop teacher-in-the-loop AI systems where educators can review and correct AI-generated responses, refining LLM performance over time. (iii) Pedagogical-aware educational agents: design agentic frameworks that align with human expert pedagogical practices.
5.3. Integration of LLM agents into Existing Educational Ecosystems
The integration of LLM agents into education presents opportunities for personalized learning, curriculum development, and project-based learning (PBL), but also raises challenges related to structured deployment, equitable access, and privacy.
@@ 5.3. Integration of LLM agents into Existing Educational Ecosystems
One major challenge is the lack of structured frameworks for integrating LLM agents into educational systems. While models like the FOKE framework (Hu and Wang, 2024) combine foundation models, knowledge graphs, and prompt engineering to provide interactive and explainable learning services, broader adoption requires scalable models that can be validated in diverse real-world educational settings. Additionally, LLMs have been explored as tools to enhance creativity and collaboration in PBL, supporting students through brainstorming, problem-solving, and project execution. However, studies indicate that their effectiveness is limited by the absence of structured guidance frameworks that help educators and students seamlessly incorporate LLM agents into PBL workflows (Zha et al., 2024). Another critical challenge is ensuring equitable access to LLM-powered educational tools, particularly in underfunded schools and institutions with limited AI infrastructure. Platforms such as AI-VERDE (Mithun et al., 2025) aim to democratize access by providing a unified LLM-as-a-platform service with built-in access control, privacy-preserving mechanisms, and budget management. However, achieving widespread adoption still depends on scalable and cost-effective deployment strategies that can support educational institutions at different resource levels.
@@
Future research should focus on developing standardized frameworks to guide the structured deployment of LLMs in personalized learning, PBL, and assessment. Expanding models like FOKE with adaptive learning strategies, multimodal content, and real-time feedback could enhance instructional effectiveness. Additionally, integrating interactive AI tutors that support student collaboration, project tracking, and contextual guidance would further improve PBL applications. Ensuring cost-effective AI deployment through cloud-based and decentralized models would make LLM-powered learning tools more accessible to a wider range of institutions. Furthermore, providing educators with AI literacy training and oversight tools is essential for responsible and effective integration.
6. Conclusion
In this survey, we provided a comprehensive review of LLM agents for education, highlighting their potential to revolutionize personalized learning, intelligent tutoring, and pedagogical automation. We introduced a task-centric taxonomy, categorizing LLM agents into Pedagogical Agents and Domain-Specific Educational Agents, and analyzed their applications across various educational domains. Furthermore, we discussed key challenges and future research directions, including ethical considerations, bias mitigation, and real-world integration. To support further advancements, we compiled essential datasets and benchmarks to facilitate research in this emerging field. As LLM agents continue to evolve, their impact on engagement, automation, and personalized learning will grow, but ensuring their effectiveness, reliability, and ethical deployment remains a key challenge. We hope this survey serves as a foundation for future research, driving innovations in AI-driven education and advancing educational equity.
Appendix A Datasets & Benchmarks

### v2

@@ 5 Challenges and Future Directions
In this section, we discuss key challenges that must be addressed to ensure the effective, reliable, and ethical deployment of LLM agents in educational settings. We focus on three critical areas: (1) privacy, bias, and fairness; (2) hallucination and overreliance; and (3) integration with existing educational ecosystems. For each of these challenges, we outline potential research directions aimed at improving the robustness, trustworthiness, and practical applicability of LLM agents in real-world education environments.
5.1 Privacy, Bias and Fairness
@@
Analysis. LLM agents process vast datasets, often containing sensitive personal information, leading to potential privacy risks. Studies highlight low technological readiness and insufficient privacy measures in educational contexts Yan et al. (2024a). Emerging research He et al. (2024a); Gan et al. (2024); Zhang et al. (2024c); Hua et al. (2024); Huo et al. (2025); Chen et al. (2025b) underscores new privacy and security concerns, emphasizing the need for stronger data protection mechanisms Huang (2023); Ismail (2025); Khan (2024). Additionally, bias in LLMs remains a pressing concern, as models trained on large datasets can inadvertently reinforce stereotypes and disparities, affecting educational fairness. Recent work calls for bias mitigation strategies to promote equitable learning experiences Adewumi et al. (2024); Aird et al. (2024); Mehrotra et al. (2024). Addressing these biases is essential to ensuring inclusive, unbiased educational outcomes.
@@
Directions. To overcome the above issues, a number of future directions can be explored: (i) Unlearning for privacy preservation: leverage advances in machine unlearning Liu et al. (2025) to enable agents to retain useful knowledge while selectively forgetting sensitive user data when required. (ii) Bias detection and mitigation: develop automated fairness-checking models that evaluate real-time content generated by LLM agents to detect biased explanations, language, or examples. (iii) Culturally adaptive LLM Agents for global education: train multilingual, culturally aware educational agents that dynamically adjust explanations based on regional learning norms, historical perspectives, and diverse curricula.
5.2 Hallucination and Overreliance
Analysis. The “hallucination” phenomenon, in which LLMs generate plausible but incorrect or nonsensical information, poses a significant challenge to their reliability in educational contexts Zhang et al. (2023). Such inaccuracies can mislead learners by presenting false information with confident and authoritative language, making errors difficult to detect and potentially leading to misconceptions da Silva et al. (2024); Jho (2024); Ho et al. (2024). For example, AI-generated content may fabricate historical events or scientific facts that students unknowingly accept as true. This risk is further amplified by overreliance, as students and educators may accept AI-generated responses without sufficient critical evaluation. Research has shown that excessive dependence on AI systems can hinder skill acquisition and reduce meaningful engagement with learning materials Milano et al. (2023); Krupp et al. (2024); Adewumi et al. (2023).
@@ 5.2 Hallucination and Overreliance
Directions. Some directions can be explored to mitigate hallucinations in LLM agents for education: (i) Self-correcting AI tutors: develop LLM agents with self-reflection capabilities Renze and Guven (2024), where models review, verify, and refine their own generated content before presenting it to students. (ii) Hybrid Human-AI feedback loops for educational content verification: develop teacher-in-the-loop AI systems where educators can review and correct AI-generated responses, refining agents performance over time. (iii) Pedagogical-aware educational agents: design agentic frameworks aligned with human pedagogical expertise to mitigate overreliance on AI-generated content.
5.3 Integration with Existing Educational Ecosystems
@@
Analysis. Although LLM agents hold great potential for automating educational practices, it is essential to consider how they can be effectively integrated into existing human-centered educational paradigms. One major challenge is the lack of structured frameworks for integrating LLM agents into educational systems. While models like the FOKE framework Hu and Wang (2024) combine foundation models, knowledge graphs, and prompt engineering to provide interactive and explainable learning services, broader adoption requires scalable models that can be validated in diverse real-world educational settings. Additionally, LLMs have been explored as tools to enhance creativity and collaboration in project-based learning (PBL), supporting students through brainstorming, problem-solving, and project execution. However, studies indicate that their effectiveness is limited by the absence of structured guidance frameworks that help educators and students seamlessly incorporate LLM agents into PBL workflows Zha et al. (2024). Another critical challenge is ensuring equitable access to LLM-powered educational tools, particularly in underfunded schools and institutions with limited AI infrastructure. Platforms such as AI-VERDE Mithun et al. (2025) aim to democratize access by providing a unified LLM-as-a-platform service with built-in access control, privacy-preserving mechanisms, and budget management. However, achieving widespread adoption still depends on scalable and cost-effective deployment strategies that can support educational institutions at different resource levels.
@@
Directions. Future research should focus on developing standardized frameworks to guide the structured deployment of LLM agents in personalized learning, PBL, and assessment. For example, expanding models like FOKE with adaptive learning strategies, multimodal content, and real-time feedback could enhance instructional effectiveness. Additionally, integrating interactive AI tutors that support student collaboration, project tracking, and contextual guidance would further improve PBL applications. To promote equitable access, develop cost-effective AI tutors through cloud-based and decentralized models would make LLM-powered learning tools more accessible to a wider range of institutions. Finally, to support meaningful integration of LLM agents into educational ecosystems, future work should move beyond task accuracy and explore more practical metrics such as learning gains, user trust, and engagement, which calls for the development of novel education-oriented benchmarks and datasets.
5.4 Multimodal LLM Agents for Education
Analysis. Despite rapid progress Bewersdorff et al. (2025); Küchemann et al. (2025), deploying multimodal LLM agents in education faces several unresolved challenges. First, integration of heterogeneous modalities such as speech, images, videos, and VR simulations is technically demanding. Models often overfit to text while underutilizing other modalities Zheng et al. (2025b). Second, cultural and contextual alignment is still limited: current agents may misinterpret classroom discourse, gestures, or culturally grounded communication patterns, reducing the accuracy and inclusiveness of feedback. Third, real-time adaptability is critical yet underexplored. For example, VR-based or speech-driven learning environments demand low-latency responses, but existing agents struggle with efficiency when reasoning across modalities Voultsiou and Moussiades (2025).
Directions. Some directions can be explored to advance the use of multimodal LLM agents in education: (i) Robust multimodal fusion: leverage contrastive pretraining and modality-specific adapters to balance contributions from text, speech, and vision, reducing overfitting to dominant modalities Nguyen et al. (2022). (ii) Context-aware personalization: incorporate localized knowledge graphs and human-in-the-loop feedback to better align agent reasoning with diverse classroom discourse and cultural norms. (iii) Latency-aware optimization: deploy lightweight distilled models and selective modality routing to ensure real-time responsiveness in VR and speech-driven learning environments.
6 Conclusion
In this survey, we presented a comprehensive review of LLM agents for education, focusing on their technical foundations and their potential to transform personalized learning, intelligent tutoring, and pedagogical automation. We proposed a task-centric taxonomy that categorizes LLM agents into Teaching Assistance and Student Support, highlighting their core capabilities such as memory augmentation, tool use, planning, and personalization. We also examined key research challenges, including ethical issues, hallucination and overreliance, and integration with existing educational ecosystems, which must be addressed to ensure reliable and ethical deployment. To support continued progress in this field, we compiled critical datasets, benchmarks, and evaluation methodologies. As LLM agents continue to evolve, their influence on education will expand. Realizing their full potential, however, will require both technical rigor and thoughtful system design. We hope this survey provides a solid foundation for future research and the development of AI-powered educational systems.
Limitations
Considering the rapid development of LLM agents for education, it is possible that some of the most recent advancements may not have been captured at the time of writing. Nevertheless, we have made every effort to ensure that all foundational and representative works are included to provide a comprehensive and accurate overview of the field. Specifically, we searched Google Scholar using keywords like “LLM agents for education”, “LLM agents for classroom simulation”, “LLM agents for adaptive learning”, and “LLM agents for knowledge tracing” covering publications from 2020 to May 2025. We included studies focusing on LLM agent-based methods but excluded those solely on LLM-based approaches.
Appendix A Domain-Specific Educational Agents
Recent research on LLM agents in education has also shown growing interest in domain-specific applications. We explore their use in science learning, language learning, and professional development, focusing on their algorithmic frameworks, agentic designs, and relevant datasets and benchmarks. In Figure 3, we present a systematic taxonomy of domain-specific educational agents.
A.1 Agent for Science Learning
An agent for science learning is an intelligent system powered by LLMs, designed to assist students in acquiring and applying scientific knowledge through personalized, interactive experiences (Yan et al., 2025b; Raihan et al., 2025; Ng et al., 2024; Brown and Cruz Castro, 2025). The significance of these agents in education lies in their ability to offer tailored feedback, enhance conceptual understanding, and promote active engagement with complex scientific ideas. In the following sections, we explore the impact of LLM agents in four key scientific disciplines: mathematicsA.1.1), physicsA.1.2), chemistryA.1.3), and biologyA.1.4), as well as their broader contributions to general scientific discoveryA.1.5).
A.1.1 Mathematics
In mathematics, LLM agents provide substantial support by helping students navigate complex problems and reinforcing their understanding of abstract concepts (Yan et al., 2024b; Xiong et al., 2024; Swan et al., 2023; Yan et al., 2024c; Wu, 2025; Mitra et al., 2024). For instance, Gou et al. (2023) introduce TORA (Tool-integrated Reasoning Agents), a framework that integrates natural language reasoning and program-based tool use to handle mathematical reasoning. MathAgent (Yan et al., 2025a) similarly proposes Mixture-of-Math-Agent framework to address multimodal error detection in real-world K-12 scenarios, and flexibly transform the visual information of different types of questions into forms that are more easily understood by LLMs (e.g., converting plane geometry images into formalized expression). Additionally, MathChat (Wu et al., 2023b) serves as a conversational mathematical problem-solving agent, which consists of a chat-based LLM agent and a tool-based user agent. Furthermore, Xiong et al. (2024) propose to use reinforcement learning from human feedback (RLHF) to further improve tool-integrated agents for mathematical problem-solving, and formulate this method as a Markov decision process, distinguishing it from the typical contextual bandit approach used in RLHF. Besides, MACM (Lei et al., 2025) discuss the limitations of LLMs in handling complex mathematical logical deduction, thus introducing a multi-agent system, which comprises three interactive agents: Thinker, Judge, and Executor.
A.1.2 Physics
In the field of physics, LLM agents help students make sense of challenging concepts and offer interactive tools to simulate physical phenomena (Pang et al., 2024; Mower and Bou-Ammar, 2025; Barman et al., 2025; Feng et al., ; Jiang and Jiang, 2024; Yan and Lee, 2024). Wang et al. (2023) introduce NEWTON, the first pipeline and benchmark to explore the physical reasoning abilities of LLMs. Furthermore, Kortemeyer (2023) describe a case study exploring if an LLM agent can pass an introductory calculus-based physics course. In addition, Physics Reasoner (Pang et al., 2024), a novel knowledge-augmented framework for physics problem-solving, leverages a comprehensive formula set and detailed checklists to ensure accuracy and completeness. It can serve as an agent consisting of three stages - problem analysis, formula retrieval, and guided reasoning. Besides, Ma et al. (2024a) describe the Scientific Generative Agent (SGA), a bilevel optimization framework designed for physical scientific discovery, and highlight the use of LLMs for generating and revising scientific hypotheses and implementing an exploit-and-explore strategy.
A.1.3 Chemistry
Chemistry education also benefits greatly from LLM agents, which can explain molecular structures, chemical reactions, and experimental processes in an engaging and interactive way (Ramos et al., 2025; M. Bran et al., 2024; Yu et al., 2024; Guo et al., 2023; Tsai et al., 2023). For example, ChemCrow (M. Bran et al., 2024) is the first LLM chemistry agent capable of autonomous planning and execution of chemical syntheses, including an insect repellent and three organocatalysts. Yu et al. (2024) further present ChemAgent, an enhanced chemistry agent improved over ChemCrow, with a focus on two essential cognitive abilities of chemistry problem-solving: reasoning and grounding. Besides, Curie (Kon et al., 2025) is an agent framework aimed at incorporating rigor into the experimentation process via three core elements: an intra-agent rigor module to boost reliability, an inter-agent rigor module to ensure systematic control, and an experiment knowledge module to improve interoperability. Recent studies have explored the capabilities of LLMs in complex chemical discovery (Yang et al., 2024d; Ruan et al., 2024; Sprueill et al., 2024; Moret et al., 2023; Jablonka et al., 2024), and their potential can be advanced by leveraging the interactivity of agent-based tool use and the flexibility of planning strategies (Song et al., 2024; Ramos et al., 2025).
A.1.4 Biology
In biology, LLM agents enhance learning by offering detailed explanations of biological processes and providing interactive experiences to explore living systems (Yan et al., 2025b; Bhattacharya et al., 2024; Sripathi et al., 2024; Zhao et al., 2025; Gao et al., 2024). For example, ProtChat (Huang et al., 2024a) is a multi-agent tool leveraging GPT-4 and Protein Language Models for seamless protein analysis automation, thus evolutionizing the complexities of protein sequence interpretation. ProtAgents (Ghafarollahi and Buehler, 2024a) is introduced as a multi-agent modeling framework that combines state-of-the-art LLMs with diverse tools to tackle protein design and analysis. It consists of a team of agents: User, Planner, Assistant, Critic, and Group Chat Manager. Besides, Shen et al. (2024) present TourSynbio-Agent, an innovative agent framework that leverages TourSynbio-7B’s protein understanding ability to perform various protein engineering tasks, such as mutation analysis, inverse folding, and visualization.
A.1.5 General Scientific Discovery
LLM agents support general scientific discovery by assisting students in data interpretation, hypothesis testing, and creative problem-solving (Yan et al., 2025b; Chen et al., 2024d; Ghafarollahi and Buehler, 2024b; Chen et al., 2025a; Schmidgall et al., 2025). These LLM agents, such as PaperQA (Lála et al., 2023), PaperQA2 (Skarlinski et al., 2024), OpenScholar (Asai et al., 2024), SciAgents (Ghafarollahi and Buehler, 2024b), TheoremExplainAgent (Ku et al., 2025), and LLM-SR (Shojaee et al., 2024), can analyze complex scientific datasets, helping students uncover patterns and trends that may not be immediately apparent. In addition, Narayanan et al. (2024) present Aviary, an extensible gymnasium for language agents for three challenging scientific tasks: manipulating DNA constructs for molecular cloning, answering research questions by accessing scientific literature, and engineering protein stability. Furthermore, both SciAgent (Ma et al., 2024c) and SciToolAgent (Chen et al., 2025a) extend to a tool-augmented scientific reasoning setting with the help of domain-specific tools. Besides, Agent Laboratory (Schmidgall et al., 2025) emerges as an agent framework that automates the research process of three phases (Literature Review, Experimentation, and Report Writing) via various LLM agents (PhD, Postdoc, ML Engineer, etc.).
A.2 Agent for Language Learning
The integration of LLM agents into language learning is revolutionizing how core competencies—reading, writing, listening, and speaking—are taught and practiced Ye et al. (2025b). These skills form the foundation of effective communication and language acquisition, and recent advancements in LLM-based agents have significantly enhanced how learners interact with and acquire these skills. Below, we introduce recent advancements in each subdomain, highlighting the role of LLM agents in enhancing pedagogical outcomes in language acquisition for students and second language (L2) speakers through engaging and adaptive approaches Huang et al. (2023b); Ye et al. (2023b, a).
A.2.1 Reading
Reading comprehension is a vital component of language learning, and LLM agents are playing an increasingly important role in enhancing students’ reading abilities. For instance, ExpertEase Mo and Hu (2024) employs a multi-agent framework to adapt documents for grade-specific audiences, simulating expert-teacher-student collaboration to enhance comprehension. AgentSimp Fang et al. (2025) tackles document-level simplification by leveraging multiple agents with distinct roles to ensure coherence and accessibility. Additionally, various LLMs Adetayo et al. (2024) have been used as academic reading companions, demonstrating improved engagement and understanding of complex qualitative texts in educational settings Chen and Leitch (2024).
A.2.2 Writing
The development of writing skills has benefited significantly from NLP tasks like explainable grammatical error correction (EXGEC) (Ye et al., 2024b, 2025a; Zou et al., 2025) and automatic essay scoring (AES) systems (Su et al., 2025). Weaver (Wang et al., 2024c), a family of LLMs fine-tuned for writing tasks Ye et al. (2023c, 2024c), outperforms generalist LLMs like GPT-4 in generating human-like narratives. Moreover, Weaver natively supports retrieval-augmented generation (RAG) and function calling, serving as a qualified foundational model for LLM agents. For interactive feedback on student essays, CAELF Hong et al. (2024) introduces a multi-agent framework that enables interactive essay feedback. By combining Teaching-Assistant agents’ evaluations with teacher-agent arbitration, students can contest grades and engage with the feedback, addressing the “black box” limitations of traditional automated scoring. Inspired by the process of human debate, Debate-to-Write Hu et al. (2025b) construct a persona-based multi-agent framework that can enable agents to collaboratively debate, discuss ideas, and form a comprehensive plan for argument writing.
A.2.3 Translation
LLM agents demonstrate remarkable advancements in both simultaneous Koshkin et al. (2024); Guo et al. (2024b) and literary translation Cheng et al. (2024); Škobo and Petričević (2023). Translation tasks benefit from LLM agents through their ability to integrate specialized tools and orchestrate multi-agent collaboration. Agent-SiMT Guo et al. (2024a) combines the decision-making capabilities of a Simultaneous Machine Translation (SiMT) policy agent with the generative power of a translation agent, achieving state-of-the-art performance in simultaneous translation by dynamically balancing reading and generation actions. For literary translation, TransAgents Wu et al. (2024a) employs a multi-agent framework to replicate the complex workflows of human translation teams, addressing cultural nuances and stylistic challenges through collaborative reasoning. This approach not only improves translation quality but also extends LLM applications to linguistically and culturally rich domains. These contributions underscore the importance of tool use and agent collaboration in advancing translation education Zheng et al. (2024).
A.2.4 Storytelling
Storytelling applications leverage LLM agents to create immersive and interactive learning experiences Simon and Muise (2022). STARie Li and Xu (2023), a peer-like embodied conversational agent, integrates multimodal tools such as speech synthesis and facial animation to scaffold children’s storytelling, fostering narrative creativity and oral communication skills Beredo and Ong (2021); Cassell (2022). StoryAgent Sohn et al. (2024) combines top-down story drafting with bottom-up asset generation to transform simple prompts into coherent, multi-modal digital narratives. By automating complex storytelling workflows Liem et al. (2023), it democratizes content creation and enhances engagement in language learning. LLaMS Zhang et al. (2025b), a multi-modal agent framework, is designed to generate multi-modal human-level stories characterized by expressiveness and consistency, incorporating the Story-Adapter module for long image sequence illustration. These systems demonstrate the potential of LLM agents to support both cognitive and creative aspects of language education for children.
A.2.5 Speaking
LLM agents are revolutionizing spoken language education by integrating reasoning and multi-agent collaboration to build adaptive dialogue systems Liu et al. (2024d); Balan et al. (2024). ELLMA-T Pan et al. (2024) employs contextual reasoning and role-playing in social VR environments to provide personalized feedback and language assessments, enabling learners to practice speaking in realistic scenarios Lim et al. (2024); Li et al. (2025b). SpokenWOZ Si et al. (2023) introduces a large-scale benchmark for task-oriented spoken dialogue, highlighting the importance of reasoning and multi-turn interaction in addressing real-world conversational challenges. FurChat Cherakara et al. (2023), an embodied conversational agent, combines verbal and non-verbal communication cues to simulate natural interactions, making it a valuable tool for improving speaking skills through immersive and realistic practice. By employing multimodal signals such as speech and gestures, SpeechAgents Zhang et al. (2024a) enhances the authenticity of dialogue simulations, capturing consistent content, natural rhythm, and rich emotional expression. Through Multi-Agent Tuning Liang et al. (2024), it optimizes LLM capabilities for large-scale simulations involving up to 25 agents, enabling applications like drama creation and audio novel generation.
A.3 Agent for Professional Development
Agents for professional development harness the capabilities of LLMs to offer scalable, adaptive, and context-aware learning experiences tailored to domain-specific needs. This section summarizes how recent studies develop agents to revolutionize professional training in fields including medicalA.3.1), computer scienceA.3.2), and law educationA.3.3).
A.3.1 Medical Education
The deployment of LLM agents in healthcare has created new opportunities for personalized, interactive, and scalable systems, with several health agents introduced Shusterman et al. (2025) such as Talk2Care Yang et al. (2024c) and openCHA Abbasian et al. (2023). Additionally, Abd-Alrazaq et al. (2023) highlight the educational potentials of LLMs in crafting personalized curricula, adaptive learning plans, and dynamic assessment tools for medical education, while concurrently addressing challenges including algorithmic bias, misinformation, and privacy issues. MEDCO Wei et al. (2024), a multi-agent system, has the capacity to replicate real-world medical training environments through agent collaboration with virtual patients, expert physicians, and radiologists, enhancing interdisciplinary learning and peer interaction. Furthermore, Abbasian et al. (2023) introduce openCHA as a personalized LLM-powered framework that integrates external resources and orchestrates multi-step problem-solving for complex healthcare queries Ye et al. (2024a), emphasizing tool use and action planning. Beyond traditional education, Steenstra et al. (2024) explore LLMs in creating fantasy narrative games for adolescent health education, demonstrating the agents’ ability to generate engaging, doctor-validated content that enhances knowledge retention through gamification. Li et al. (2024b) present Agent Hospital, a simulation environment where LLM-driven agents evolve through autonomous interactions, demonstrating significant improvements in medical reasoning and performance on benchmarks like MedQA Jin et al. (2021) after treating thousands of simulated patients. Collectively, these investigations highlight the versatility of LLM agents within medical education, demonstrating their abilities in reasoning, collaboration, tool integration, and adaptive learning to effectively address a broad spectrum of educational and clinical challenges Karabacak and Margetis (2023); Tian et al. (2024b); Ullah et al. (2024).
A.3.2 Computer Science Education
An agent for computer science (CS) education greatly enhances learning by providing personalized guidance on coding, debugging, and understanding CS principles Ma et al. (2024b); Lee and Song (2024); Kosar et al. (2024); Liu et al. (2024b). For example, CodeAgent Zhang et al. (2024b) serves as an LLM agent framework for repo-level code generation, incorporating external tools such as WebSearch and DocSearch. Recent studies have demonstrated the potential of agent-based code generation systems such as ToolCoder Ding et al. (2025), SWE-agent Yang et al. (2025), AgentCoder Huang et al. (2023a), and MapCoder Islam et al. (2024), which can significantly enhance students’ coding efficiency Jin et al. (2024a); Wang et al. (2025a); Frankford et al. (2024). Furthermore, Bassner et al. (2024) introduce Iris, an LLM-driven virtual tutor designed to offer personalized, context-aware assistance to CS students within the interactive learning platform Artemis. Besides, Chen et al. (2024a) propose Learning-by-Teaching (LBT) as an effective pedagogical strategy for CS education, and leverage the advantages of LLM agents (e.g., contextual conversation & learning from demonstrations).
A.3.3 Law Education
LLM agents leverage pre-trained legal knowledge, interactive capabilities, and reasoning skills to support law education through judicial interpretation, moot court simulation, and case analysis Chen et al. (2024b); Nelson (2024); Lai et al. (2024); Yuan et al. (2024). However, evaluations from LawBench Fei et al. (2023) and LegalBench Guha et al. (2023) reveal that LLMs struggle with legal knowledge application and judicial aid. LegalAgentBench (Li et al., 2024a) further highlights their limitations in multi-hop reasoning and defense statement writing, showing that LLM agents require significant improvements to effectively assist in complex legal tasks. Despite these challenges, LLM agents are emerging as valuable tools for moot court simulations, a crucial component of legal reasoning and advocacy training. DeliLaw Xie et al. (2024a) enhances law education by integrating legal and case retrieval modules, enabling students to practice legal research, case analysis, statutory interpretation, and mock consultations. LawLuo Sun et al. (2024) applies a multi-agent framework with retrieval-augmented generation to simulate multi-turn legal consultations, improving personalization and ambiguity handling. Similarly, AgentCourt Chen et al. (2024b) and AgentsCourt He et al. (2024b) simulate courtroom interactions and judicial decision-making, providing a realistic training ground for law students. AgentsBench Jiang and Yang (2024) extends this by offering multi-agent legal reasoning and case analysis, further advancing AI-driven legal education.
Appendix B Datasets & Benchmarks